# Certifying Stability of Reinforcement Learning Policies using Generalized Lyapunov Functions

### v2

@@ 6 Joint Synthesis of Stable Neural Policies and Certificates
So far, we addressed the certification of stability for fixed pre-trained control policies. In this section, we consider joint synthesis of neural controllers and Lyapunov certificates and employ formal verification inspired by Chang et al. (2019); Dai et al. (2021); Wu et al. (2023); Yang et al. (2024). We demonstrate that our generalized Lyapunov approach integrates naturally into this setting by replacing the standard one-step decrease condition with the proposed multi-step, weighted formulation in (9).
@@
In general, precise characterizations of $\mathcal{R}$ are challenging to obtain and one instead looks for suitable approximations. Here, we seek to jointly learn a policy $\boldsymbol{\pi}(\mathbf{x};\boldsymbol{\phi})$ and a certificate function $V(\mathbf{x};\boldsymbol{\theta}_{1})$ such that the closed-loop system $\mathbf{x}_{k+1}=\mathbf{f}(\mathbf{x}_{k},\boldsymbol{\pi}_{\phi}(\mathbf{x}_{k}))$ is provably asymptotically stable and obtain at the same time an inner approximation $\mathcal{S}$ of the ROA $\mathcal{R}$. Yang et al. (2024) formalized it as a constrained optimization that maximizes the volume of a Lyapunov sublevel set $\mathcal{S}=\{\mathbf{x}\in\mathbb{R}^{n}:V(\mathbf{x})\leq\rho\}$ under Lyapunov constraints
@@
$F(\mathbf{x}_{k}):=\frac{1}{M}\sum_{i=1}^{M}\sigma_{i}(\mathbf{x}_{k})V(\mathbf{x}_{k+i};\boldsymbol{\theta}_{1})-(1-\bar{\alpha})V(\mathbf{x}_{k};\boldsymbol{\theta}_{1})\leq 0,\;\;\sigma_{i}(\mathbf{x}_{k})\geq\underline{\sigma}>0,$ (18)
@@
Training Formulation. Following (Yang et al., 2024), we reformulate the problem (17) into a learning objective by sampling states $\mathbf{x}_{k}\in\mathcal{X}$ and penalizing violations of the Lyapunov conditions using soft constraints. For each $\mathbf{x}_{k}\in\mathcal{D}$, we simulate $M$ forward steps under the closed-loop dynamics and compute the generalized Lyapunov residual $F(\mathbf{x}_{k})$ as defined in (18).
@@
To expand the certifiable region, Yang et al. (2024) proposed a surrogate region loss $\mathcal{L}_{\text{region}}:=\sum_{j=1}^{N}\text{ReLU}\left(V(\mathbf{x}_{j};\boldsymbol{\theta}_{1})/\rho-1\right)$, where the candidate states $\mathbf{x}_{j}$ are obtained via random boundary sampling or projected gradient descent (PGD) to minimize $V(\cdot;\boldsymbol{\theta}_{1})$. PGD is also used for falsification by maximizing stability violations, generating additional training states for $\mathcal{D}$.
@@
Although the theoretical framework (Theorem 4.2) supports trainable weights $\sigma_{i}(\mathbf{x}_{k})$, including neural network parameterizations (see Section 5), certifying stability under jointly learned controllers, Lyapunov functions, and weights remains computationally challenging with existing tools (Wang et al., 2021; Gao et al., 2013). In practice, we fix $\sigma_{i}$ during training. For small values of $M$, we select the best-performing weight configuration via a simple grid search.
Verification Formulation. After training, we verify that the generalized decrease condition holds within $\mathcal{X}$ using the $\alpha$-$\beta$-CROWN verifier (Wang et al., 2021). We formally certify Lyapunov stability of the origin if the following holds for all $\mathbf{x}_{k}\in\mathcal{X}$:
@@
As Theorem 6.2 implies, if (20) holds for all $\mathbf{x}_{k}\in\mathcal{X}$, then $\mathcal{S}$ is an inner approximation of ${\cal R}$.
Evaluation. We evaluate our generalized Lyapunov synthesis approach on three systems presented in (Yang et al., 2024): inverted pendulum, path tracking, and 2D quadrotor (with a 6D state space). Detailed system dynamics and neural network architectures used are provided in Appendix F.
Figures 8 and 8 show the certified stability regions $\mathcal{S}$ for different horizon lengths $M$. We use fixed step weights $(\sigma_{1},\sigma_{2},\dots)$ selected via grid search: for the inverted pendulum, $(0.4,1.6)$ for $M{=}2$ and $(0.3,1.5,1.2)$ for $M{=}3$; for path tracking, $(0.4,1.6)$ and $(1.2,1.2,0.6)$, respectively. As shown in the figures, multi-step training and verification yields consistently larger certifiable ROAs.
@@
Table 2 presents the quantitative results, including the volume of $\mathcal{S}$ and verification time. The volume is estimated via Monte Carlo integration: we sample $n=10^{6}$ points in $\mathcal{X}$ and compute the fraction satisfying $V(\mathbf{x})\leq\rho$, averaged over 10 trials and scaled by the total domain volume. While our generalized stability formulation leads to consistently larger certified regions, it incurs longer formal verification time due to the need to bound multiple intermediate states and their weighted combinations, leading to more complex bound computations and more branching decisions.
Table 2: Certified region volume (left) and verification time (right) under different $M$-step Lyapunov training.
@@
System $M=1$ $M=2$ $M=3$
Inverted Pendulum 11.54 22.42 39.87
Path Tracking 8.53 19.61 36.26
2D quadrotor 2209.75 3824.19 5657.86
@@ 7 Conclusion, Limitations, and Future Work
This work opens several promising directions for future research. A key challenge is to enable formal verification of learned weights, certificate functions, and neural controllers using certified tools. Additionally, it is important to investigate the scalability of our framework to higher-dimensional systems. Finally, we are interested in further exploring the connection between stability and optimality, particularly in using generalized Lyapunov conditions to guide reward design and policy refinement.
@@
Appendix B Construction of Classical Lyapunov Function
This section shows that a generalized $M$-step Lyapunov function with uniformly positive step weights can be used to construct a classical one-step Lyapunov function. Specifically, if the closed-loop system $\mathbf{x}_{k+1}=\mathbf{f}(\mathbf{x}_{k},\boldsymbol{\pi}(\mathbf{x}_{k}))$ admits a generalized Lyapunov function $V:{\cal X}\to\mathbb{R}_{\geq 0}$ in the sense of Definition 4.1, then under mild assumptions on the weights, one can explicitly construct a classical Lyapunov function. This result provides a converse-type implication: the existence of a generalized Lyapunov function enables the explicit construction of a classical Lyapunov function.
Assumption B.1 (Uniformly positive weights).
There exists a constant $\underline{\sigma}>0$ such that
$\sigma_{i}(\mathbf{x})\;\geq\;\underline{\sigma},\qquad\forall\mathbf{x}\in{\cal X},\;i=1,\dots,M.$ (30)
Proposition B.2 (Construction of Classical Lyapunov Function).
Let $V$ satisfy the generalized Lyapunov condition in Definition 4.1 and assume the step weights satisfy Assumption B.1. Choose
$0<\beta<\min\left\{1,\left(\frac{\underline{\sigma}}{M}\right)^{1/(M-1)}\right\},$ (31)
and define the sliding-window sum
$W(\mathbf{x}_{k}):=\sum_{i=0}^{M-1}\beta^{i}V(\mathbf{x}_{k+i}).$ (32)
Then $W$ is a classical Lyapunov function:
$W(\mathbf{0}_{n})=0,\quad W(\mathbf{x})>0$ $\quad\forall\mathbf{x}\neq\mathbf{0}_{n},$ (33)
$W\big{(}\mathbf{f}(\mathbf{x},\boldsymbol{\pi}(\mathbf{x}))\big{)}-W(\mathbf{x})<0$ $\quad\forall\mathbf{x}\neq\mathbf{0}_{n},$ (34)
Therefore, the origin is asymptotically stable on ${\cal X}$.
@@ Appendix B Construction of Classical Lyapunov Function
By Definition 4.1 and Assumption B.1, we have
$V(\mathbf{x}_{k+i})\leq\frac{M}{\underline{\sigma}}V(\mathbf{x}_{k}),\qquad\forall i=1,\dots,M.$ (35)
Using the definition of $W$ in (32) and shifting the index:
$W(\mathbf{x}_{k+1})=\sum_{i=0}^{M-1}\beta^{i}V(\mathbf{x}_{k+1+i})=\beta^{-1}\left(W(\mathbf{x}_{k})-V(\mathbf{x}_{k})+\beta^{M}V(\mathbf{x}_{k+M})\right).$ (36)
Substituting the bound (35) for $V(\mathbf{x}_{k+M})$, we obtain:
$W(\mathbf{x}_{k+1})-W(\mathbf{x}_{k})\leq\left(\beta^{M-1}\frac{M}{\underline{\sigma}}-1\right)V(\mathbf{x}_{k}).$ (37)
By the choice of $\beta$ in (31), the bracketed term is strictly negative, so $W(\mathbf{x}_{k+1})<W(\mathbf{x}_{k})$ for all $\mathbf{x}_{k}\neq\mathbf{0}_{n}$.
To verify positive definiteness, note that $V(\mathbf{x}_{k})>0$ for all $\mathbf{x}_{k}\neq\mathbf{0}_{n}$ implies each term in the sum (32) is nonnegative, and at least one is strictly positive. Therefore, $W(\mathbf{x}_{k})>0$ for all $\mathbf{x}_{k}\neq\mathbf{0}_{n}$, and $W(\mathbf{0}_{n})=\sum_{i=0}^{M-1}\beta^{i}V(\mathbf{0}_{n})=0$, since $V(\mathbf{0}_{n})=0$ by assumption.
Thus, $W$ is a classical Lyapunov function certifying asymptotic stability of the origin. ∎
Hence a uniformly positive multi-step certificate automatically yields a classical Lyapunov function; one can view Proposition B.2 as an alternative to the proof of Theorem 4.2 under additional mild assumptions.
Remark B.3 (Zero weights).
When Assumption B.1 is dropped, the sliding–window sum (32) may not decrease, so the construction in Proposition B.2 may fail. Nevertheless, Theorem 4.2 shows that a generalized Lyapunov function with vanishing weights still certifies asymptotic stability. By the discrete-time converse Lyapunov theorem, some classical Lyapunov function must therefore exist, but it may not admit a closed-form like (32). Finding an explicit formula in the zero-weight case remains an open problem and is left for future work.
@@
Following the approach in [Postoyan et al., 2017], we introduce an auxiliary quadratic function:
@@ D.1 Network Architectures
We use feedforward neural networks with LeakyReLU activations for both the residual network $\varphi(\cdot,\boldsymbol{\theta}_{1})$ and step-weighting network $\sigma(\cdot,\boldsymbol{\theta}_{2})$. The architectures are summarized in Table 3. The residual network is initialized using Kaiming initialization [He et al., 2015], and all biases are initialized to zero.
@@ D.3 Compute Resources
Stability. Let $\epsilon>0$ be given. Since $V$ is continuous and positive definite, there exists $\delta_{0}>0$ such that:
$\|\mathbf{x}\|<\delta_{0}\quad\Rightarrow\quad V(\mathbf{x})<\epsilon.$ (56)
Define the constant:
$c:=\frac{2(1-\bar{\alpha})}{\underline{\sigma}}>0,$ (57)
and choose:
$\delta:=\min\left\{\delta_{0},\frac{\epsilon}{c}\right\}.$ (58)
Suppose $\mathbf{x}_{k}\in\mathcal{S}$ and $\|\mathbf{x}_{k}\|<\delta$. Then by (56) and (58), we have:
$V(\mathbf{x}_{k})<\frac{\epsilon}{c}.$ (59)
Applying the generalized Lyapunov condition (55) and using $\sigma_{1},\sigma_{2}\geq\underline{\sigma}$, we obtain:
$\underline{\sigma}\left(V(\mathbf{x}_{k+1})+V(\mathbf{x}_{k+2})\right)\leq\sigma_{1}V(\mathbf{x}_{k+1})+\sigma_{2}V(\mathbf{x}_{k+2})\leq 2(1-\bar{\alpha})V(\mathbf{x}_{k}),$ (60)
which implies that for $i=1,2$:
$V(\mathbf{x}_{k+i})\leq cV(\mathbf{x}_{k})<\epsilon,$ (61)
where we used (57) and (59). Then by (56), we conclude:
$\|\mathbf{x}_{k+1}\|<\epsilon,\quad\|\mathbf{x}_{k+2}\|<\epsilon.$ (62)
Moreover, due to the strict decrease enforced by (55), it is not possible for both $V(\mathbf{x}_{k+1})\geq V(\mathbf{x}_{k})$ and $V(\mathbf{x}_{k+2})\geq V(\mathbf{x}_{k})$ to hold simultaneously. Therefore, at least one of the two future states must satisfy $V(\mathbf{x}_{k+i})<V(\mathbf{x}_{k})$ for $i\in\{1,2\}$, indicating progress toward the origin. In particular, this ensures that within at most two steps, the trajectory returns to a strictly smaller sublevel set of $\mathcal{S}$, allowing the generalized Lyapunov condition to be re-applied.

### v4

@@ 6 Joint Synthesis of Stable Neural Policies and Certificates
So far, we addressed the certification of stability for fixed pre-trained control policies. In this section, we consider joint synthesis of neural controllers and Lyapunov certificates and employ formal verification inspired by Chang2019NeuralLC; dai_2021_lyapunov; wu2023neural; yang2024lyapunovstable. Compared to prior methods, our generalized Lyapunov framework naturally extends this setting by replacing the standard one-step decrease condition with the proposed multi-step, weighted formulation in (9).
@@
In general, precise characterizations of $\mathcal{R}$ are challenging to obtain and one instead looks for suitable approximations. Here, we seek to jointly learn a policy $\boldsymbol{\pi}(\mathbf{x};\boldsymbol{\phi})$ and a certificate function $V(\mathbf{x};\boldsymbol{\theta}_{1})$ such that the closed-loop system $\mathbf{x}_{k+1}=\mathbf{f}(\mathbf{x}_{k},\boldsymbol{\pi}_{\phi}(\mathbf{x}_{k}))$ is provably asymptotically stable and obtain at the same time an inner approximation $\mathcal{S}$ of the ROA $\mathcal{R}$. yang2024lyapunovstable formalized it as a constrained optimization that maximizes the volume of a Lyapunov sublevel set $\mathcal{S}=\{\mathbf{x}\in\mathbb{R}^{n}:V(\mathbf{x})\leq\rho\}$ under Lyapunov constraints
@@
$F(\mathbf{x}_{k}):=\frac{1}{M}\biggl(\sum_{i=1}^{M}\sigma_{i}(\mathbf{x}_{k})V(\mathbf{x}_{k+i};\boldsymbol{\theta}_{1})\biggr)-(1-\bar{\alpha})V(\mathbf{x}_{k};\boldsymbol{\theta}_{1})\leq 0,\;\;\sigma_{i}(\mathbf{x}_{k})\geq\underline{\sigma}>0,$ (18)
@@
Training Formulation. Following (yang2024lyapunovstable), we reformulate the problem (17) into a learning objective by sampling states $\mathbf{x}_{k}\in\mathcal{X}$ and penalizing violations of the Lyapunov conditions using soft constraints. For each $\mathbf{x}_{k}\in\mathcal{D}$, we simulate $M$ forward steps under the closed-loop dynamics and compute the generalized Lyapunov residual $F(\mathbf{x}_{k})$ as defined in (18).
@@
To expand the certifiable region, yang2024lyapunovstable proposed a surrogate region loss $\mathcal{L}_{\text{region}}:=\sum_{j=1}^{N}\text{ReLU}\left(V(\mathbf{x}_{j};\boldsymbol{\theta}_{1})/\rho-1\right)$, where the candidate states $\mathbf{x}_{j}$ are obtained via random boundary sampling or projected gradient descent (PGD) to minimize $V(\cdot;\boldsymbol{\theta}_{1})$. PGD is also used for falsification by maximizing stability violations, generating additional training states for $\mathcal{D}$.
@@
Although the theoretical framework (Theorem 4.2) supports trainable weights $\sigma_{i}(\mathbf{x}_{k})$, including neural network parameterizations (see Section 5), certifying stability under jointly learned controllers, Lyapunov functions, and weights remains computationally challenging with existing tools (wang2021beta; gao2013dreal). In practice, we fix $\sigma_{i}$ during training. For small values of $M$, we select the best-performing weight configuration via a simple grid search.
Verification Formulation. After training, we verify that the generalized decrease condition holds within $\mathcal{X}$ using the $\alpha$-$\beta$-CROWN verifier (wang2021beta). We formally certify Lyapunov stability of the origin if the following holds for all $\mathbf{x}_{k}\in\mathcal{X}$:
@@
[…]
Evaluation. We evaluate our generalized Lyapunov synthesis approach on three systems presented in (yang2024lyapunovstable): inverted pendulum, path tracking, and 2D quadrotor (with a 6D state space). Detailed system dynamics and neural network architectures used are provided in Appendix F.
Figures 8 and 8 show the certified stability regions $\mathcal{S}$ for different horizon lengths $M$. We use fixed step weights $(\sigma_{1},\sigma_{2},\dots)$ selected via grid search: for the inverted pendulum, $(0.4,1.6)$ for M=2M{=}2 and $(0.3,1.5,1.2)$ for M=3M{=}3; for path tracking, $(0.4,1.6)$ and $(1.2,1.2,0.6)$, respectively. As shown in the figures, multi-step training and verification yields consistently larger certifiable ROAs.
@@
Table 3 reports the quantitative results, including the volume $(\mathcal{S})$ and the verification time. The volume is estimated via Monte Carlo integration: we sample $n=10^{6}$ points in $\mathcal{X}$ and compute the fraction satisfying $V(\mathbf{x})\leq\rho$, averaged across trials and scaled by the total domain volume. To assess robustness, we repeat each experiment 10 times with different random seeds and report the mean and standard deviation. The results confirm that our generalized stability formulation consistently certifies larger ROA volumes compared to the classical one-step Lyapunov approach. This improvement comes at the cost of longer verification times, since bounding multiple intermediate states and their weighted combinations increases the complexity of bound propagation and branching decisions.
Table 3: Certified region volume (left) and verification time (right) under different $M$-step Lyapunov training. Abbreviations: I.P. = Inverted Pendulum, P.T. = Path Tracking, 2.Q. = 2D Quadrotor.
@@
$M=1$ $M=2$ $M=3$
11.7 $\pm$ 0.5 21.5 $\pm$ 0.8 39.2 $\pm$ 1.3
8.6 $\pm$ 0.5 19.5 $\pm$ 1.0 36.7 $\pm$ 1.3
2209.5 $\pm$ 72.6 3858.7 $\pm$ 112.4 5628.6 $\pm$ 185.9
@@ 7 Conclusion, Limitations, and Future Work
This work opens up several promising directions for future research. One key avenue is to develop partial-state stability certification, providing guarantees for task-relevant components such as torso height or velocity in humanoid locomotion rather than for the full system state, which can improve both scalability and interpretability. A second direction is to extend the generalized Lyapunov formulation to certify robust stability under bounded perturbations and stochastic disturbances, taking inspiration from input-to-state and robust Lyapunov theory. Another important direction is to relax the requirement of knowing the equilibrium state, allowing stability analysis when the equilibrium is unknown or non-isolated, as often occurs in reinforcement learning, by inferring equilibria from data or from the reward structure. Finally, investigating the connection between generalized Lyapunov conditions and optimality objectives may yield new insights for reward design and policy refinement in reinforcement learning.
Acknowledgments
We gratefully acknowledge support from ONR Award N00014-23-1-2353 and NSF Award CCF-2112665 (TILOS).
@@ Acknowledgments
Appendix B Construction of a Classical Lyapunov Function from a Generalized One
This appendix shows that when the generalized $M$-step Lyapunov decrease in Definition 4.1 holds with state-independent step weights, one can explicitly construct a classical one-step Lyapunov function, following a similar idea to classical constructions that convert multi-step (non-monotone) decrease conditions into a one-step Lyapunov decrease amir_2008_non_montonic_LF. In contrast, when the step weights depend on the state, obtaining an explicit closed-form classical Lyapunov function from the generalized certificate appears non-trivial; while asymptotic stability still follows from Theorem 4.2, deriving a simple one-step Lyapunov function construction in the state-dependent case is left for future work.
Assumption B.1 (Constant step weights).
The step weights in Definition 4.1 are constant, i.e., there exist $\sigma_{1},\dots,\sigma_{M}\geq 0$ (independent of $\mathbf{x}$) such that for all $\mathbf{x}\in\mathcal{X}\setminus\{\mathbf{0}\}$,
$\frac{1}{M}\Big(\sum_{i=1}^{M}\sigma_{i}V(\mathbf{x}_{i})\Big)-V(\mathbf{x})<0,\;\;\frac{1}{M}\sum_{i=1}^{M}\sigma_{i}\geq 1,$ (30)
where $\mathbf{x}_{0}=\mathbf{x}$ and $\mathbf{x}_{i+1}=\mathbf{f}(\mathbf{x}_{i},\boldsymbol{\pi}(\mathbf{x}_{i}))$ for $i=0,\dots,M-1$.
Lemma B.2 (Explicit construction of a classical Lyapunov function).
Suppose $V:\mathcal{X}\to\mathbb{R}_{\geq 0}$ satisfies Definition 4.1 and Assumption B.1. Without loss of generality, assume
$\sum_{i=1}^{M}\sigma_{i}=M.$ (31)
Define coefficients
$a_{j}:=\frac{1}{M}\sum_{i=j+1}^{M}\sigma_{i},\qquad j=0,\dots,M-1,$ (32)
and the function
$W(\mathbf{x})~:=~\sum_{j=0}^{M-1}a_{j}\,V\!\left(\mathbf{f}^{\,j}(\mathbf{x})\right),$ (33)
where $\mathbf{f}^{\,j}$ denotes $j$-fold composition of the closed-loop map $\mathbf{x}\mapsto\mathbf{f}(\mathbf{x},\boldsymbol{\pi}(\mathbf{x}))$. Then $W$ is a classical Lyapunov function:
$W(\mathbf{0})=0,\qquad W(\mathbf{x})>0\ \ \forall\mathbf{x}\neq\mathbf{0},$ (34)
and it decreases strictly in one step:
$W\!\left(\mathbf{f}(\mathbf{x},\boldsymbol{\pi}(\mathbf{x}))\right)-W(\mathbf{x})<0\ \ \forall\mathbf{x}\neq\mathbf{0}.$ (35)
@@ Appendix B Construction of a Classical Lyapunov Function from a Generalized One
By (31) and (32), we have $a_{0}=1$ and $a_{j}\geq 0$ for all $j$. Hence
$W(\mathbf{x})\geq V(\mathbf{x}),$ (36)
which implies (34). If $V$ is radially unbounded, then so is $W$ by (36).
Let $\{\mathbf{x}_{k}\}_{k\geq 0}$ be any closed-loop trajectory and define $V_{k}:=V(\mathbf{x}_{k})$. Using (33),
$W(\mathbf{x}_{k+1})-W(\mathbf{x}_{k})$ $=\sum_{j=0}^{M-1}a_{j}V_{k+1+j}-\sum_{j=0}^{M-1}a_{j}V_{k+j}$ (37)
$=-a_{0}V_{k}+\sum_{j=1}^{M-1}(a_{j-1}-a_{j})V_{k+j}+a_{M-1}V_{k+M}.$
By (32), $a_{j-1}-a_{j}=\sigma_{j}/M$ for $j=1,\dots,M-1$ and $a_{M-1}=\sigma_{M}/M$, and by (31) we have $a_{0}=1$. Substituting into (37) gives
$W(\mathbf{x}_{k+1})-W(\mathbf{x}_{k})=\frac{1}{M}\sum_{i=1}^{M}\sigma_{i}V_{k+i}-V_{k}.$ (38)
The right-hand side of (38) is strictly negative for all $\mathbf{x}_{k}\neq\mathbf{0}$ by (30), proving (35). ∎
Scaling the weights.
If $\sum_{i=1}^{M}\sigma_{i}>M$, define $\tilde{\sigma}_{i}:=\frac{M}{\sum_{j=1}^{M}\sigma_{j}}\sigma_{i}$. Then $\tilde{\sigma}_{i}\geq 0$ and $\sum_{i=1}^{M}\tilde{\sigma}_{i}=M$. Since $\frac{M}{\sum_{j=1}^{M}\sigma_{j}}\in(0,1)$, the left-hand side of (30) is scaled down while $V(\mathbf{x})$ is unchanged, so the strict inequality in (30) remains valid. Hence (31) is without loss of generality.
@@
Following the approach in [Romain_2017_TAC], we introduce an auxiliary quadratic function:
@@ D.1 Network Architectures
We use feedforward neural networks with LeakyReLU activations for both the residual network $\varphi(\cdot,\boldsymbol{\theta}_{1})$ and step-weighting network $\sigma(\cdot,\boldsymbol{\theta}_{2})$. The architectures are summarized in Table 4. The residual network is initialized using Kaiming initialization [he2015delving], and all biases are initialized to zero.
@@ D.3 Compute Resources
Stability. We follow the classical $\varepsilon$$\delta$ Lyapunov proof in [khalil1996nonlinear, Thm. 3.3]. Fix an arbitrary radius $\varepsilon>0$ and define
$m_{\varepsilon}\;:=\;\inf_{\|x\|=\varepsilon}V(x).$ (57)
Because the $\varepsilon$‑sphere $\mathcal{E}=\{x\in\mathbb{R}^{n}:\|x\|=\varepsilon\}$ is compact and $V$ is continuous, the infimum is achieved on $\mathcal{E}$. Positive‑definiteness of $V$ further implies $m_{\varepsilon}>0$.
Define
$c\;:=\;\frac{2(1-\bar{\alpha})}{\underline{\sigma}}\;>\;0.$
Because $V$ is continuous at the origin and $V(0)=0$, for every threshold $\eta>0$ there exists a radius $\delta(\eta)>0$ such that
$\|x\|<\delta(\eta)\;\;\Longrightarrow\;\;V(x)<\eta.$
Choosing $\eta=m_{\varepsilon}/c$ and setting $\delta=\delta(m_{\varepsilon}/c)$, gives
$\|x\|<\delta\;\;\Longrightarrow\;\;V(x)<\frac{m_{\varepsilon}}{c}.$ (58)
Now, pick any index $k$ with $x_{k}\in\mathcal{S}$ and $\|x_{k}\|<\delta$. From (58), we have the bound
$V(x_{k})\;<\;\frac{m_{\varepsilon}}{c}.$ (59)
Using the two–step Lyapunov condition (56) and the fact that $\sigma_{1},\sigma_{2}\geq\underline{\sigma}$, we obtain
$\underline{\sigma}\!\left(V(x_{k+1})+V(x_{k+2})\right)\;\leq\;2(1-\bar{\alpha})\,V(x_{k})\;\;\Longrightarrow\;\;V(x_{k+i})\;\leq\;c\,V(x_{k}),\quad i=1,2.$ (60)
Combining (60) with (59) gives
$V(x_{k+1})<m_{\varepsilon},\qquad V(x_{k+2})<m_{\varepsilon}.$
By definition of $m_{\varepsilon}$, $V(x)<m_{\varepsilon}$ implies $\|x\|<\varepsilon$; hence
$\|x_{k+1}\|<\varepsilon,\qquad\|x_{k+2}\|<\varepsilon.$ (61)
Equation (61) shows that, starting from any state inside the $\delta$‑ball, the trajectory remains inside the prescribed $\varepsilon$‑ball.
Moreover, due to the strict decrease enforced by (56), it is not possible for both $V(\mathbf{x}_{k+1})\geq V(\mathbf{x}_{k})$ and $V(\mathbf{x}_{k+2})\geq V(\mathbf{x}_{k})$ to hold simultaneously. Therefore, at least one of the two future states must satisfy $V(\mathbf{x}_{k+i})<V(\mathbf{x}_{k})$ for $i\in\{1,2\}$, hence the trajectory enters a strictly smaller sublevel set of $V$ within two steps, allowing the generalized Lyapunov condition to be re-applied.